# Exponentiations of ultrafilters

### v3

@@ 2   Exponentiations of ultrafilters
$(2)$ If $p=q$, both equivalences are trivial.
If $n^{p}=n^{q}$ it means that for all $A\subseteq\mathbb{N}$, if $B=\{a\in\mathbb{N}\mid n^{a}\in A\}$ then $B\in p\Leftrightarrow B\in q$, which shows that if we let $P_{n}=\{a\in\mathbb{N}\mid\exists b\in\mathbb{N}\ a=b^{n}\}$ and $log_{n}:P_{n}\rightarrow\mathbb{N}$ be the $n$-base logarithm and $\widehat{log_{n}}$ be its extension to ultrafilters containing $P_{n}$ then $p=\widehat{log_{n}}\left(n^{p}\right)=\widehat{log_{n}}\left(n^{q}\right)=q$.
The other equivalence can be proven similarly.
$(3)$ If $n=m$, both equivalences are trivial.
If $n^{p}=m^{p}$, first we observe that $n^{\mathbb{N}}=\{n^{k}\mid k\in\mathbb{N}\}\in n^{p}$ and $m^{\mathbb{N}}=\{m^{k}\mid k\in\mathbb{N}\}\in m^{p}$. Therefore, if $n^{p}=m^{p}$ we have that $n^{\mathbb{N}}\cap m^{\mathbb{N}}\in n^{p}$. Hence $n^{\mathbb{N}}\cap m^{\mathbb{N}}\neq\emptyset$, which forces $n^{a}=m$ for some $a\in\mathbb{Q}$, so $n^{p}=n^{ap}$ from $(1)$, which from $(2)$ forces $p=ap$, which forces $a=1$, so $n=m$.
The other equivalence is similar.
$(4)$ First, let us show that $n^{p}\odot n^{q}=n^{p\oplus q}$: let $A\subseteq\mathbb{N}$. Then
$A\in n^{p}\odot n^{q}\Leftrightarrow\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid ab\in A\}\in n^{q}\}\in n^{p}\Leftrightarrow$
$\{d\in\mathbb{N}\mid n^{d}\in\{a\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c}\in\{b\in\mathbb{N}\mid ab\in A\}\in q\}\}\}\in p\Leftrightarrow$
$\{d\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c}\cdot n^{d}\in A\}\in q\}\in p\Leftrightarrow\{d\in\mathbb{N}\mid\{c\in\mathbb{N}\mid n^{c+d}\in A\}\in q\}\in p\Leftrightarrow$
$\{e\in\mathbb{N}\mid n^{e}\in A\}\in p\oplus q\Leftrightarrow A\in n^{p\oplus q}.$
For the other claims: let $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}$ and let $P\in r$. Let $p,q$ be nonprincipal. Then a direct computation shows that
@@
and we conclude as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
$(5)$ Let $A\subseteq\mathbb{N}$. Then
$A\in E_{1}(p,q\odot r)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\odot r\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid n^{ab}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(n^{a}\right)^{b}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(n^{a}\right)^{b}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{c\in\mathbb{N}\mid\{d\in\mathbb{N}\mid c^{d}\in A\}\in r\}\in E_{1}\left(p,q\right)\Leftrightarrow$
$A\in E_{1}\left(E_{1}\left(p,q\right),r\right).$
$(6)$ This is analogous to $(5)$. Let $A\subseteq\mathbb{N}$. Then
$A\in E_{2}\left(p,E_{2}(q,r)\right)\Leftrightarrow\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in E_{2}(q,r)\}\in p\Leftrightarrow$
$\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid\left(b^{a}\right)^{n}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{n\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid b^{an}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
$\{c\in\mathbb{N}\mid\{b\in\mathbb{N}\mid b^{c}\in A\}\in r\}\in p\odot q.\qed$
Example 2.5.
Their definition should make evident that $E_{1},E_{2}$ are ultrafilters operations with relationships with exponential configurations.
@@
Proceeding inductively like above, we construct a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ with the property that for all $j<i$ one has that
$F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A.$
For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j<i\}\subseteq A$.
Moreover, by choosing $p$ with additional properties, one could get more complicated configurations. For example, by taking $p$ such that each set in $p$ contains arbitrarily long arithmetic progressions, we can take all sets $F_{i}$ to be arithmetic progressions of length $i$ in $B$.
Example 2.7.
If we argue as above with $E_{2}$ instead of $E_{1}$, we get analogous results with the order relationship between $i$ and $j$ switched. So, in the same notations, we get a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ with the property that for all $j<i$ one has that
$F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A.$
For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j>i\}\subseteq A$.
@@
To conclude this section, we want to show that the operations $E_{1},E_{2}$ are related to exponential triples. This should not come as a surprise as, even if using a different language, Di Nasso and Ragosta in [2] to prove the exponential Schur theorem employed ultrafilters $u$ of the form $u=2^{q}$ where $q$ had the form $q=2^{p}\odot p$ for some $p$ witness of van der Waerden Theorem. In [4], the authors proved the following result, which allows to generalize the observation of Di Nasso and Ragosta::
Theorem 2.8.
Let $A$ be a multiplicatively piecewise syndetic set. Then $A$ contains an exponential triple.
As an immediate corollary, the following holds.
Corollary 2.9.
Let $\mathcal{U}\in\overline{K(\beta\mathbb{N},\odot)}$. Then every $A\in\mathcal{U}$ contains an exponential triple.
Notice that, by Corollary 2.9, it is simple to prove that several configurations are partition regular by using [1, Lemma 2.1]. To recall it, we need to introduce the following notation: for $p\in\beta\mathbb{N}$, $f\left(x_{1},\dots,x_{n}\right):\mathbb{N}^{n}\rightarrow\mathbb{N}$, we write $p\models f\left(x_{1},\dots,x_{n}\right)=0$ if for all $A\in\mathcal{U}$ there exists $a_{1},\dots,a_{n}\in A$ such that $f\left(a_{1},\dots,a_{n}\right)=0$.
Lemma 2.10.
[1, Lemma 2.1] Let $p\in\beta\mathbb{N}$, and let $f\left(x_{1},\dots,x_{n}\right),g\left(y_{1},\dots,y_{m}\right)$ be two functions such that $p\models\left(f_{1}\left(x_{1},\dots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\dots,y_{m}\right)=0\right)$. Then $p\models\left(f_{1}\left(x_{1},\dots,x_{n}\right)=0\right)\wedge\left(g\left(y_{1},\dots,y_{m}\right)=0\right)\wedge\left(x_{1}=y_{1}\right)$.
As said, the above result allows us to recover the partition regularity of many equations and configurations. In fact, pick $p\in\overline{K(\beta\mathbb{N},\odot)}$ multiplicatively idempotent. Then:
[…]
  • as $p$ is idempotent, $p\models y_{1}\dot{}\dots\dot{y}_{m}=z$;
@@
and so on, as we can always substitute a variable $x$ with a polynomial $P\left(y_{1},\dots,y_{n}\right)$ whenever $p\models x=P\left(y_{1},\dots,y_{n}\right)$. We refer to [1, Theorem 2.11] for a long list of such admissible substitutions.
Corollary 2.9 can be generalized also in another direction. Let us first fix a definition.
Definition 2.11.
We let
$EVdW=\{p\in\beta\mathbb{N}\mid\forall A\in p,\forall k\in\mathbb{N}\,\exists\,x,y\in A\ \text{s.t.}\ x^{y},\dots,x^{\left(y^{k}\right)}\in A\}.$
@@
For all $\mathcal{U}\in\overline{K(\beta\mathbb{N},\odot)}$, for all $n\in\mathbb{N}_{>1}$, $n^{\mathcal{U}}\in EVdW$.
Actually, Theorem 2.12 can be seen as a very particular case of a much more general phenomenon involving $E_{1},E_{2}$.
Theorem 2.13.
Let $p,q\in\beta\mathbb{N}\setminus\{0\}$. Then:
  1. 1.
    If $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\in EVdW$. In particular, $E_{1}\left(p,q\right)$ witnesses the partition regularity of exponential triples;
  2. 2.
    If $q\in EVdW$ then $E_{2}(p,q)\in EVdW$. In particular, $E_{2}(p,q)$ witnesses the partition regularity of exponential triples.
@@
(1) Let $k\in\mathbb{N}$ and let $A\in E_{1}(p,q)$. We will show that $A$ contains a configuration of the form $x,y,x^{y},\dots,x^{\left(y^{k}\right)}$.
By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in q\}\in p$. Let $n\in I$ and $J_{n}=\{m\in\mathbb{N}\mid n^{m}\in A\}$. Take $a,b\in J_{n}$ with $a,b,an^{b},\dots,an^{kb}\in J_{n}$, whose existence is ensured by Corollary 2.9 since, as $J_{n}\in q\in\overline{K(\beta\mathbb{N},\odot)}$, $J_{n}$ is multiplicatively piecewise syndetic set. Therefore,
$n^{a},n^{b},n^{an^{b}}=\left(n^{a}\right)^{\left(n^{b}\right)},\dots,n^{an^{kb}}=\left(n^{a}\right)^{\left(\left(n^{b}\right)^{k}\right)}\in A,$
and we conclude by letting $x=n^{a},y=n^{b}$.
(2) Let $q\in EVdW$, let $k\in\mathbb{N}$ be given and let $A\in E_{2}(p,q)$. By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid m^{n}\in A\}\in q\}\in p$. Let $n\in I$ and let $J_{n}=\{m\in\mathbb{N}\mid m^{n}\in A\}$. Take $a,b\in J_{n}$ with $a,b,a^{\left(b^{n}\right)},\dots,a^{\left(b^{kn}\right)}\in J_{n}$. Hence
$a^{n},b^{n},\left(a^{\left(b^{n}\right)}\right)^{n}=a^{nb^{n}}=\left(a^{n}\right)^{\left(b^{n}\right)},\dots,\left(a^{\left(b^{kn}\right)}\right)^{n}=a^{nb^{kn}}=\left(a^{n}\right)^{\left(\left(b^{n}\right)^{k}\right)}\in A,$
and we conclude by letting $x=a^{n},y=b^{n}$.∎
As a consequence of [6, Theorem 6], if $Sc_{+}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, then $Sc_{+}\cap EVdW=\emptyset$; in particular,
$\overline{K(\beta\mathbb{N},\odot)}\cap EVdW=\emptyset,\overline{E(\oplus)}\cap EVdW=\emptyset,$
as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq Sc_{+}$. Then, in particular, by Theorem 2.12 we also deduce that for all $n>1$, for all $p\in\overline{K(\beta\mathbb{N},\odot)}$, $n^{p}\notin\overline{K(\beta\mathbb{N},\odot)}$.
@@
As it is nowadays well-known, idempotents are often key ingredients in the search for infinite partition regular structures, the most famous example being the relationship between idempotent ultrafilters with respect to the sum and Hindman Theorem (see [5, Chapter 5]). As Examples 2.6, 2.7 suggest, having introduced the binary operations $E_{1},E_{2}$, it makes sense to study if they admit idempotents:

### v4

@@ 2   Exponentiations of ultrafilters
$(2)$ $\Rightarrow$)⇒ ) Consider the function $\log_{a}:\mathbb{R}^{+}\rightarrow\mathbb{R}$. A straightforward computation shows that $\widehat{\log_{a}}\left(a^{p}\right)=p$, hence our hypothesis entails that $p=\widehat{\log_{a}}\left(b^{q}\right)$. Now we observe that, for any $A\subseteq\mathbb{R}^{+}$,
$A\in\widehat{\log_{a}}\left(b^{q}\right)\Leftrightarrow(\log_{a})^{-1}(A)\in\left(b^{q}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in(\log_{a})^{-1}(A)\}\in q\}\in b\Leftrightarrow$
$\{y\in\mathbb{R}^{+}\mid b^{y}\in(\log_{a})^{-1}(A)\}\in q\Leftrightarrow\{y\in\mathbb{R}^{+}\mid\log_{a}\left(b^{y}\right)\in A\}\in q\Leftrightarrow$
$\{y\in\mathbb{R}^{+}\mid y\log_{a}(b)\in A\}\in q\Leftrightarrow A\in\left(\log_{a}(b)\right)q.$
Hence $p=\left(\log_{a}(b)\right)q$, as desired.
$\Leftarrow$)⇐ ) Just observe that $b^{q}=b^{\left(\log_{a}(b)\right)p}$, and $b^{\left(\log_{a}(b)\right)p}=E_{1}\left(b^{\left(\log_{a}(b)\right)},p\right)$ by $(1)$. We conclude as $b^{\left(\log_{a}(b)\right)}=a$.
$(3)$ $\Rightarrow$)⇒ ) This time, consider the function $f(x):=\sqrt[a]{x}:\mathbb{R}^{+}\rightarrow\mathbb{R}$. Again, a straightforward computation shows that $\widehat{f}\left(p^{a}\right)=p$, hence $p=\widehat{f}\left(q^{b}\right)$. For any $A\subseteq\mathbb{R}^{+}$ we have that
$A\in\widehat{f}\left(q^{b}\right)\Leftrightarrow(f)^{-1}(A)\in\left(q^{b}\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in f^{-1}(A)\}\in q\}\in b\Leftrightarrow$
$\{y\in\mathbb{R}^{+}\mid y^{b}\in f^{-1}(A)\}\in q\Leftrightarrow\{x\in\mathbb{R}^{+}\mid f\left(x^{b}\right)\in A\}\in q\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid x^{\frac{b}{a}}\in A\}\in q\Leftrightarrow A\in q^{\frac{b}{a}}.$
$\Leftarrow$)⇐ ) By $(1)$ we have that $q^{b}=E_{1}\left(p^{\frac{a}{b}},b\right)=E_{1}\left(p,\frac{a}{b}\cdot b\right)=p^{a}$.
$(4)$ Let $A\subseteq\mathbb{R}^{+}$. Then
$A\in a^{p}\odot a^{q}\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid xy\in A\}\in a^{q}\}\in a^{p}\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid a^{t}\in\{y\in\mathbb{R}^{+}\mid xy\in A\}\}\in q\}\in a^{p}\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid x\cdot a^{t}\in A\}\in q\}\in a^{p}\Leftrightarrow\{z\in\mathbb{R}^{+}\mid a^{z}\in\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid x\cdot a^{t}\in A\}\in q\}\in p\Leftrightarrow$
$\{z\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid a^{t+z}\in A\}\in q\}\in p\Leftrightarrow A\in a^{p\oplus q}.$
$(5)$ Let $A\subseteq\mathbb{R}^{+}$. Then
$A\in E_{1}(p,q\odot r)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid x^{y}\in A\}\in q\odot r\}\in p\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid x^{tz}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid\left(x^{t}\right)^{z}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
$\{y\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid s^{z}\in A\}\in r\}\in E_{1}\left(p,q\right)\Leftrightarrow A\in E_{1}\left(E_{1}\left(p,q\right),r\right).$
$(6)$ This is analogous to $(5)$. Let $A\subseteq\mathbb{R}^{+}$. Then
$A\in E_{2}\left(p,E_{2}(q,r)\right)\Leftrightarrow\{x\in\mathbb{R}^{+}\mid\{y\in\mathbb{R}^{+}\mid y^{x}\in A\}\in E_{2}(q,r)\}\in p\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid\left(z^{t}\right)^{x}\in A\}\in r\}\in q\}\in p\Leftrightarrow$
$\{x\in\mathbb{R}^{+}\mid\{t\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid z^{tx}\in A\}\in r\}\in q\}\in p\Leftrightarrow\{s\in\mathbb{R}^{+}\mid\{z\in\mathbb{R}^{+}\mid z^{s}\in A\}\in r\}\in p\odot q.\qed$
Corollary 2.11.
Let $p,q,\beta\mathbb{N}\setminus{1}$, $a,b\in\mathbb{N}\setminus\{1\}$. Then
  1. 1.
    $a^{p}=a^{q}\Leftrightarrow p=q$;
  2. 2.
    $p^{a}=p^{b}\Leftrightarrow a=b$.
Proof.
(1) This follows from Proposition 2.10.(2).
(2) This follows from Proposition 2.10.(3), as $p^{\frac{a}{b}}=p$ if and only if $\frac{a}{b}=1$. ∎
Example 2.12.
Proposition 2.10 might again suggest that many arithmetic properties of the exponentiations of natural numbers hold also for $E_{1},E_{2}$. For example, Proposition 2.10.(4) generalizes the usual rule $a^{b}\cdot a^{c}=a^{bc}$ to the case where $b,c$ are ultrafilters. However, this same rule cannot be generalized to the case where $a$ is an ultrafilter, as we show in the next proposition.
Proposition 2.13.
Let $p\in\beta\mathbb{N}\setminus\{\mathbb{N}\}$, and let $a,b\in\mathbb{N}$. Then $p^{a}\odot p^{b}\neq p^{a+b}$.
Proof.
Assume, by contrast, that $p^{a}\odot p^{b}=p^{a+b}$.
Let $F:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps a number $n$ to its largest prime divisor. Observe that, for all $p_{1},p_{2}\in\beta\mathbb{N}$, $\hat{\Omega}(E_{2}(p_{1},p_{2}))=p_{1}\odot\hat{\Omega}(p_{2})$ and $\hat{\Omega}(p_{1}\odot p_{2})=\hat{\Omega}(p_{1})\oplus\hat{\Omega}(p_{2})$. We consider two cases.
Case 1: $\hat{F}(p)=i\in\mathbb{N}$. In this case, consider the function $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ that maps a number $n$ to the sum of the exponents in its prime factorization. Necessarily, $\hat{\Omega}(p)\notin\mathbb{N}$; in fact, if by contrast $\hat{\Omega}(p)=j\in\mathbb{N}$, we would have that $p\in\overline{S}$, where
$S=\{n\in\mathbb{N}\mid F(n)=i,\Omega(n)=j\},$
which is a contradiction as $p$ is nonprincipal and $S$ is finite. Hence $\hat{\Omega}(p)\notin\mathbb{N}$. As $p^{a}\odot p^{b}=p^{a+b}$, we have that
$\hat{\Omega}\left(p^{a}\odot p^{b}\right)=\hat{\Omega}\left(p^{a}\right)\oplus\hat{\Omega}\left(p^{b}\right)=a\hat{\Omega}(p)\oplus b\hat{\Omega},$
whilst
$\hat{\Omega}\left(p^{a+b}\right)=(a+b)\hat{\Omega}(p).$
We conclude as $a\hat{\Omega}(p)\oplus b\hat{\Omega}(p)\neq(a+b)\hat{\Omega}(p)$ by Theorem 2.6.
Case 2: $\hat{F}(p)\notin\mathbb{N}$. Let $G:\mathbb{N}\rightarrow\mathbb{N}$ be the function $n\rightarrow\max\{m\in\mathbb{N}\mid F(n)^{m}\ \text{divides}\ n\}$, and let $H(n):=F(n)^{G(n)}$. As $\hat{F}(p)$ is nonprincipal, also $\hat{H}(p)$ is nonprincipal. And
$\hat{H}\left(p^{a}\odot p^{b}\right)=b\hat{H}(p)\neq(a+b)\hat{H}(p)=\hat{H}\left(p^{a+b}\right),$
again by Theorem 2.6. ∎
Remark 2.14.
The above results prove that the equalities $E_{1}(r,p)\odot E_{1}(r,q)=E_{1}(r,p\oplus q),E_{2}(p,r)\odot E_{2}(q,r)=E_{2}(p\oplus q,r)$ does not hold when $r$ is principal. They do not need to hold even when $r$ is nonprincipal. E.g., assume that $r$ is nonprincipal and $P=\{n\in\mathbb{N}\mid n\ \text{is a prime}\}\in r$. A direct computation shows that
@@
and we conclude that $E_{1}(r,p)\odot E_{1}(r,q)\neq E_{1}(r,p\oplus q)$ and $E_{2}(p,r)\odot E_{2}(q,r)\neq E_{2}(p\oplus q,r)$ as $I_{1}\cap J=I_{2}\cap J=\emptyset.$
Their definition should make evident that $E_{1},E_{2}$ are ultrafilters operations that can be used to study the partition regularity of certain exponential configurations.
@@
Proceeding inductively like above, it is possible to construct a sequence $\langle F_{i}\rangle_{i=1}^{\infty}$ of finite subsets of $B_{0}$ such that, for all $j<i$, $F_{j}^{F_{i}}=\{m^{n}\mid m\in F_{j},n\in F_{i}\}\in A$. For example, if $F_{i}=\{a_{i}\}$ for all $i$, we obtain an infinite subset $X=\{a_{i}\mid i\in\mathbb{N}\}$ of $B_{0}$ such that $\{a_{j}^{a_{i}}\mid a_{i},a_{j}\in X,j<i\}\subseteq A$. Arguing as above with $E_{2}$ instead of $E_{1}$, we could deduce an analogous results with the order relationship between $i$ and $j$ switched.
@@
We now want to show that the operations $E_{1},E_{2}$ are related to the partition regularity of exponential triples, as well as of other exponential configurations. This should not come as a surprise: the main results in [2] can be restated in terminology introduced in Section 2.1 as follows:
Theorem 2.16 ([2, Theorem 2.1, restated]).
If $p$ is a Van der Waerden’s witness then $2^{p}\odot p\models\{x,y,2^{x}y\}$.
From this fact, in [2, Corollary 2.3] the authors deduced the exponential Schur’s Theorem by the $\log_{2}$-partition method mentioned in the introduction. Their proof actually shows the following fact:
Theorem 2.17 ([2, Theorem 2.3, restated]).
If $p$ is Van der Waerden’s witness then $2^{2^{p}\odot p}\models\{x,y,x^{y}\}$.
As it is well-known (see e.g. [5, Theorem 14.5]), any $p\in\overline{K(\beta\mathbb{N},\odot)}$ is a Van der Waerden’s witness; moreover, as $p\in\overline{K(\beta\mathbb{N},\odot)}$, we have that $2^{p}\odot p\in\overline{K(\beta\mathbb{N},\odot)}$. This simple observation, joint with Theorem 2.17 and Lemma 2.4, allows to prove easily that several configurations are partition regular. In fact, let $p\in\overline{K(\beta\mathbb{N},\odot)}$, and let $q=2^{p}$. Then:
@@
  • by furtherly assuming that $p=p\odot p$, one gets that $q\models x^{\log_{2}(y)}=z$, hence e.g. $q\models x_{1}^{x_{2}\log_{2}x_{3}}=y_{1}\cdot y_{2}^{y_{3}}$;
  • if $p$ is also combinatorially rich, then $p\models x=y+z^{2}$ (see e.g. [1, Theorem 2.11], hence $q\models x=yt^{\log_{2}(t)}$,
and so on, as we can always substitute a variable $x$ with a polynomial $P\left(y_{1},\ldots,y_{n}\right)$ whenever $q\models x=P\left(y_{1},\ldots,y_{n}\right)$. We refer to [1, Theorem 2.11] for a long list of such admissible substitutions.
To conclude this section we study some properties of the set of exponential Schur’s witnesses, namely
$\mathrm{ESW}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists x,y\in A\ x^{y}\in A\}.$
The set $\mathrm{ESW}$ has some nice closure property. To prove them, we will need the following strengthening of Di Nasso and Ragosta’s results, whose proofs are almost identical to those of the results in [2]. We add them here for completeness.
@@
Let $n\in\mathbb{N}$. If $p$ is a Van der Waerden’s witness then $n^{p}\odot p\models\{x,y,n^{x}y\}$.
@@
Let $A\in n^{p}\odot p$. By definition, this means that $A^{\prime}=\{a\in\mathbb{N}\mid A/n^{a}\in p\}\in p$, where $A/n^{a}=\{m\in\mathbb{N}\mid n^{a}m\in A\}$. Pick $a\in A^{\prime}$. As $A^{\prime}\cap A/n^{a}\in p$, there are $b,c,b+n^{a}c\in A/n^{a}\cap A^{\prime}$. In particular, $A/n^{b},A/n^{b+n^{a}c}\in p$. Let $d\in A/n^{b}\cap A/{n^{b+n^{a}c}}$. Letting $x=n^{a}c,y=n^{b}d$, we have that $x\in A$ as $c\in A/n^{a}$, $y\in A$ as $d\in A/n^{b}$, and $n^{x}y=n^{n^{a}c}n^{b}d=n^{b+n^{a}c}d\in A$ as $d\in A/n^{b+n^{a}c}$.∎
Corollary 2.19.
Let $n\in\mathbb{N}$. If $p$ is a Van der Waerden’s witness then $n^{n^{p}\odot p}\in\mathrm{ESW}$.
Proof.
Let $A\in n^{n^{p}\odot p}$. Let $I_{A}\in\{m\mid n^{m}\in A\}\in n^{p}\odot p$. Let $a,b\in I_{A}$ be such that $n^{a}b\in I_{A}$. Then $x:=n^{b},y:=n^{a}$ are elements of $A$ such that $x^{y}=\left(n^{b}\right)^{\left(n^{a}\right)}=n^{bn^{a}}\in A$. ∎
Our first result is that Corollary 2.19 can be easily generalized as follows.
Proposition 2.20.
Let $q\in\beta\mathbb{N}\setminus\{0,1\}$. If $p$ is a Van der Waerden’s witness then $E_{1}\left(q,E_{1}(q,p)\odot p\right)\in\mathrm{ESW}$.
Proof.
Let $A\in E_{1}\left(q,E_{1}(q,p)\odot p\right)$. By definition,
$\underbrace{\{n\in\mathbb{N}\mid\overbrace{\{m\in\mathbb{N}\mid\{a\in\mathbb{N}\mid\{b\in\mathbb{N}\mid n^{m^{a}b}\in A\}\in p\}\in p\}}^{J}\in q\}}_{I}\in q.$
If $n\in I\cap J$ then $A\in n^{n^{p}\odot p}$, and we conclude by Corollary 2.19.∎
Similarly, a more restrictive condition on $p$ allows to prove the following:
Proposition 2.21.
Let $p,q\in\beta\mathbb{N}\setminus\{0,1\}$. Assume that, for all $n>1$, $p\models\{x,y,xn^{y}\}$. Then $E_{1}\left(q,p\right)\in\mathrm{ESW}$.
Proof.
Let $A\in E_{1}(q,p)$. By definition, $I=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A\}\in p\}\in q$. Let $n\in I$ and $J_{n}=\{m\in\mathbb{N}\mid n^{m}\in A\}$. Take $a,b\in J_{n}$ with $a,b,an^{b}\in J_{n}$. Therefore
$n^{a},n^{b},n^{an^{b}}=\left(n^{a}\right)^{\left(n^{b}\right)}\in A,$
and we conclude by letting $x=n^{a},y=n^{b}$. ∎
By [5, Theorem 4.40], any $q\in\overline{K(\beta\mathbb{N},\odot)}$ consists solely of multiplicatively piecewise syndetic sets; hence, by Theorem 2.5, any such $q$ satisfies that, for all $n>1$, $q\models\{x,y,xn^{y}\}$. Therefore, Theorem 2.21 has the following immediate corollary:
Corollary 2.22.
Let $q\in\overline{K(\beta\mathbb{N},\odot)}$. Then for all $p\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $E_{1}\left(p,q\right)\in\mathrm{ESW}$.
Remark 2.23.
If $Sc_{+}=\{p\in\beta\mathbb{N}\mid\forall A\in p\ \exists a,b\in A\ a+b\in A\}$, Theorem 2.7 tells us that $Sc_{+}\cap\mathrm{ESW}=\emptyset$; in particular, if $E(\oplus)=\{p\in\beta\mathbb{N}\mid p\oplus p=p\}$ then
$\overline{K(\beta\mathbb{N},\odot)}\cap\mathrm{ESW}=\emptyset,\overline{E(\oplus)}\cap\mathrm{ESW}=\emptyset,$
as $\overline{K(\beta\mathbb{N},\odot)}\cup\overline{E(\oplus)}\subseteq Sc_{+}$. By Theorem 2.21.(1), we get that if $q\in\overline{K(\beta\mathbb{N},\odot)}$ then $E_{1}\left(p,q\right)\notin\overline{K(\beta\mathbb{N},\odot)}$; in particular, for $p,q\in\overline{K(\beta\mathbb{N},\odot)}$, $E_{1}(p,q)\notin\overline{K(\beta\mathbb{N},\odot)}$: the smallest bilateral multiplicative ideal of $\beta\mathbb{N}$ is very much not closed under the $E_{1}$ exponentiation.
@@
Having introduced two natural binary operations on $\beta\mathbb{N}$, it makes sense to ask if they admit idempotents, defined as follows.